OpenStudy (anonymous):

Marry Ann building a fence around the outer edge of a rectangular garden plot that measures 25 feet by 45 feet. She will set the posts 5 feet apart. How many posts will she need?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

24 9 for the sides of 45 and three for the sides of 25 feet. why omly three and not five? because we have posts at the corners already

OpenStudy (anonymous):

Wouldn't it be 28 posts?

OpenStudy (anonymous):

see my reason above

OpenStudy (anonymous):

Wouldn't you do the perimeter then divide by 5?

OpenStudy (anonymous):

yes. but you would be putting two in each corner.

OpenStudy (anonymous):

25/5 = 5 45/5 = 9 but now look 5 is for example the upper side and down side so the right side will have only 7 so as the left side 5 + 5 + 7 + 7 = 24

OpenStudy (anonymous):

I need to draw a diagram that represents this problem.

OpenStudy (anonymous):